§ 1) Find V2 3 V 5 V 2) Find I -6 A 2 A
Answer #1

applying kul to Loop-A Vz = 2+5= 7V. Loop-3, Vis -3- V3 = -3-7 Vi s-10v Looprie Neo H + Vi : 4-10 5-6V Vg3-6v A applying KEL